5. Find the general solution to the differential equation y'' + y= sin2 x


1
Expert's answer
2021-10-01T16:25:34-0400

Solution.


"y''+y=\sin^2(x)"

1) Solve linear equation with constant coefficients

"y''+y=0."

Сompose a characteristic equation "\lambda^2+1=0."

From here "\lambda=i," or "\lambda=-i."

So, we have


"y=C_1\sin x+C_2\cos x."

2) Find particular solution by the method of undefined coefficients. Such as "\sin^2(x)=\frac{1-\cos{2x}}{2}," particular solution for the right side equal to the sum of particular solutions for "\frac{1}{2}" and "-\frac{\cos(2x)}{2}."

Solution for "\frac{1}{2}" is:

"y=A." Then "y''=0." Substitute in original equation "A=\frac{1}{2}." So,


"y=\frac{1}{2}."


Solution for "-\frac{\cos(2x)}{2}" is:

"y=B\sin(2x)+D\cos(2x)." Then "y''=-4B\sin(2x)-4D\cos(2x)." Substitute in original equation "-6B\sin(2x)-6D\cos(2x)=-\cos(2x)." From here "B=0, D=\frac{1}{6}." So,

"y=\frac{\cos(2x)}{6}."

3)


"y=C_1\sin x+C_2\cos x+\frac{1}{2}+\frac{\cos(2x)}{6}."

Answer. "y=C_1\sin x+C_2\cos x+\frac{1}{2}+\frac{\cos(2x)}{6}."